Please help me solve this: Clifford "TI" Harris went on a weekend trip. He averaged 35mi/h in the city and 60mi/h on the highway. If the 150mi trip took 3hrs, how much of the time was he driving in the city.
Answer by mducky2(62)


Here are some equations that might help us.
Speed * Time = Distance
Time = Distance/Speed
Speed = Distance/Time

For ease of labeling, let's label time, distance, and speed with variables:
T = time
D = distance
S = speed

Now let's figure out what information the problem gives us.
speed in city = Sc = 35 mph
speed on highway = Sh = 60 mph
total time = Ttotal = 3 hours
total distance = Dtotal = 150 miles

It looks like we can use the first equation to help us. The speed times the time on the city equals the distance in the city. The speed times the time on the highway equals the distance on the highway. Adding the two together should give us the total distance:
Speed * Time = Distance
Sc*Tc = Dc
Sh*Th = Dh
ScTc+ShTh = Dtotal
35Tc + 60Th = 150

Since we know the total time, we can solve for the time in the highway in terms of the time in the city:
Tc + Th = 3
Th = 3 - Tc

Plugging back into the original equation:
35tc + 60(3 - tc) = 150
35tc + 180 - 60tc) = 150
180 - 25Tc = 150
180 - 150 = 25Tc
30 = 25Tc
Tc = 30/25
Tc = 6/5 hrs
Tc = 1 hr and 12 minutes

Clifford "TI" Harris spent 1 hr and 12 minutes driving in the city.